OpenStudy (solomonzelman):

Why is it so that x^0, no matter what the x is, is always equal to 1? (is it just rule, or there is a logic behind it?)

hartnn (hartnn):

except for x=0 ofcourse.

hartnn (hartnn):

you know whats an identity for multiplication ?

hartnn (hartnn):

like identity for addition is 0 anything +0 = anything.

OpenStudy (solomonzelman):

I don't think so, 0^0....actually not sure, hold.....

OpenStudy (solomonzelman):

yep 0^0 is 1, check on google calc.

hartnn (hartnn):

google calc is incorrect. http://www.wolframalpha.com/input/?i=0%5E0&dataset=

OpenStudy (solomonzelman):

Weird isn't it, but...... Anyway, is there a logic behind that or it is just an explainable rule.

OpenStudy (solomonzelman):

First time I see google calc is incorrect... let me check your....

OpenStudy (solomonzelman):

it says indeterminate, is that no solution?

hartnn (hartnn):

x^2 is x multiplied by itself twice x^1 is x multiplied by 1 x^0 means 'x' is multiplied 0 times, (not even once) to 1 so, 1* (nothing to multiply) = 1 yes, its one of the indeterminate forms, others include 0/0, infinity.infinity

OpenStudy (solomonzelman):

get it!

OpenStudy (solomonzelman):

I always argued w/ my elementary teacher that 0/0 =0 ....

hartnn (hartnn):

i meant inf/inf

OpenStudy (solomonzelman):

still, get it no solush...

OpenStudy (solomonzelman):

And anyway, for x^0=1 for any x value beside 0, why is that true?

hartnn (hartnn):

yup, but no solution sounds weird, there's no equation to solve. indeterminate form seems accurate

OpenStudy (solomonzelman):

Yes

hartnn (hartnn):

because 1 is the identity of multiplication!

hartnn (hartnn):

when you take any number 'x' and do not multiply it (0 times), you are only remain with 1

OpenStudy (amistre64):

1*(undefined) is undefined .... not 1

OpenStudy (solomonzelman):

you aren't multiplying by itself or a/t it is just to the 0 power. it is nothing then

OpenStudy (amistre64):

x^0 is a property of the exponential function. exp(a+b) = exp(a) * exp(b) exp(n-n) = exp(n) * exp(-n)\[e^n*e^{-n}=\frac{e^n}{e^n}=1\]

OpenStudy (solomonzelman):

70%

OpenStudy (solomonzelman):

70% get it....

OpenStudy (solomonzelman):

Yes I get it....!

OpenStudy (solomonzelman):

Thank you! i though there was no logic, but this is just b/c it is 0.... hard to explain and prove that I get it, but I get it. Thank You again, PRO!!!!

OpenStudy (amistre64):

good luck

OpenStudy (solomonzelman):

thank you!